### v2

# short-time Fourier transform iSTFT inverse short-time Fourier transform

STFT
short-time Fourier transform
iSTFT
inverse short-time Fourier transform
DNN
deep neural network
PESQ
Perceptual Evaluation of Speech Quality
POLQA
perceptual objectve listening quality analysis
WPE
weighted prediction error
PSD
power spectral density
RIR
room impulse response
SNR
signal-to-noise ratio
LSTM
long short-term memory
POLQA
Perceptual Objective Listening Quality Analysis
SDR
signal-to-distortion ratio
ESTOI
extended short-term objective intelligibility
ELR
early-to-late reverberation ratio
TCN
temporal convolutional network
RLS
recursive least squares
ASR
automatic speech recognition
HA
hearing aid
CI
cochlear implant
MAC
multiply-and-accumulate
GAN
generative adversarial network
T-F
time-frequency
SDE
stochastic differential equation
ODE
ordinary differential equation
DRR
direct to reverberant ratio
LSD
log spectral distance
SI-SDR
scale-invariant signal to distortion ratio
MOS
mean opinion score
MAP
maximum a posteriori
SDE
stochastic differential equation
ODE
ordinary differential equation
DPS
diffusion posterior sampling
VAE
variational auto-encoding
NMF
non-negative matrix factorization
WPE
weighted-prediction error
EM
expectation-maximization
Unsupervised Blind Joint Dereverberation and Room Acoustics Estimation with Diffusion Models
[…]
Abstract
This paper presents an unsupervised method for single-channel blind dereverberation and room impulse response (RIR) estimation, called BUDDy. The algorithm is rooted in Bayesian posterior sampling: it combines a likelihood model enforcing fidelity to the reverberant measurement, and an anechoic speech prior implemented by an unconditional diffusion model. We design a parametric filter representing the RIR, with exponential decay for each frequency subband. Room acoustics estimation and speech dereverberation are jointly carried out, as the filter parameters are iteratively estimated and the speech utterance refined along the reverse diffusion trajectory. In a blind scenario where the RIR is unknown, BUDDy successfully performs speech dereverberation in various acoustic scenarios, significantly outperforming other blind unsupervised baselines. Unlike supervised methods, which often struggle to generalize, BUDDy seamlessly adapts to different acoustic conditions. This paper extends our previous work by offering new experimental results and insights into the algorithm’s versatility. We demonstrate the robustness of our proposed method to new acoustic and speaker conditions, as well as its adaptability to high-resolution singing voice dereverberation, using both instrumental metrics and subjective listening evaluation. We study BUDDy’s performance for RIR estimation and observe it surpasses a state-of-the-art supervised DNN-based estimator on mismatched acoustic conditions. Finally, we investigate the sensitivity of informed dereverberation methods to RIR estimation errors, thereby motivating the joint acoustic estimation and dereverberation design. Audio examples and code can be found online.111uhh.de/sp-inf-buddy
Index Terms:
Acoustics, reverberation, speech enhancement.
I Introduction
Reverberation is a natural phenomenon caused by acoustic waves propagating in a space and reflecting off various surfaces, such as walls, ceilings, floors, and objects within the environment. Reverberation and particularly late reflections often degrade speech intelligibility and quality for normal listeners, and even more severely so for hearing-impaired listeners [1]. Therefore, many communication devices now include a dereverberation algorithm, which aims to recover the anechoic component of speech. This paper considers the case in which recordings from only one microphone are available, which is more challenging than a multi-channel case [2].
Traditional dereverberation algorithms operate in the time, spectral, or cepstral domain [3], leveraging statistical assumptions about the anechoic and reverberant signals [4] as well as properties of the reverberation signal model [5]. Two scenarios are considered for dereverberation, depending on the knowledge of the room acoustics represented by the room impulse response (RIR). Some methods tackle informed scenarios, where the RIR is known [6, 5], whereas other approaches consider blind scenarios where the RIR is unknown [7, 8, 9, 10, 11]. Informed dereverberation is naturally an easier task than blind dereverberation. However, knowing the RIR does not guarantee obtaining a stable and causal inverse filter in the single-channel case, since real-world RIRs are mixed-phase systems [12]. Using multiple microphones helps resolve this issue to some extent [2], but informed dereverberation methods generally exhibit other weaknesses such as a lack of robustness to RIR estimation errors [13]. Additionally, most scenarios in real-life applications are (at least partially) blind, as the RIR is either not measured beforehand, or only valid for a specific acoustic setting.
Data-driven approaches rely less on distributional assumptions than statistical methods but instead directly learn the signal properties and structures from data [14]. Most of these methods are based on supervised learning, where models are trained using paired data. Each input is associated with a corresponding target output, allowing the model to learn a mapping from inputs to outputs. For dereverberation, this typically involves using pairs of anechoic and reverberant speech, where the latter is often produced by convolving anechoic speech signals with RIRs. Supervised predictive models are particularly popular for blind dereverberation: these range from time-frequency masking [15] and mapping[16] to algorithms operating on the cepstrum [17] or directly on the waveform [18, 19].
Generative modeling is another paradigm gaining a lot of interest in audio restoration tasks [20], including dereverberation. Generative models for speech dereverberation learn a parameterization of the posterior distribution of clean speech conditioned on reverberant speech. Diffusion models in particular [21, 22, 23] have been extensively investigated for such conditional generation task, leading to the introduction of diffusion-based blind supervised dereverberation algorithms [24, 25]. Still, the generalization ability of supervised approaches is limited by their design.
In contrast, unsupervised methods operate without paired data, relying solely on patterns learned from anechoic speech signals. These approaches have been getting little visibility but boast interesting properties such as improved robustness to unseen acoustic conditions without the need for retraining. An unsupervised method for informed single-channel dereverberation based on diffusion models was proposed in our prior work [26]. That approach is based on Bayesian diffusion posterior sampling (DPS) [27], combining a diffusion-based anechoic speech prior and a Gaussian likelihood model for state-of-the-art informed dereverberation. However, as shown in this work, such an informed algorithm is sensitive to even small RIR estimation errors, rendering it impractical in real-life scenarios.
Related works in other signal processing domains have already considered blind inverse problems through the lens of posterior sampling with diffusion priors. For image deblurring, Chung et al. [28] propose to use an additional diffusion process dedicated to estimating the deblurring kernel, while Laroche et al. [29] adapts an expectation-maximization algorithm using a denoising regularization of the blurring kernel, and Sanghvi et al. [30] dedicates a non-blind solver to estimate a deblurred image at each diffusion step. For speech denoising, Nortier et al. [31] combine a noise model based on non-negative matrix factorization with a clean speech diffusion prior. Moliner et al. [32] address the problem of blind bandwidth extension by leveraging a diffusion prior and iteratively optimizing a parametric lowpass filter operator. Recent works adapt denoising diffusion restoration models (DDRM) [33] for singing voice dereverberation [34, 35], using an initialization provided by the weighted-prediction error (WPE) algorithm [7].
For speech dereverberation, a first generative model based on traditional Gaussian mixtures was proposed in [36]. Other works learn an anechoic speech prior via variational auto-encoding (VAE): the VAE-NMF method [37] models reverberation via non-negative matrix factorization and estimates its parameters with a Monte-Carlo method; the RVAE-EM model [38] adopts a maximum a posteriori perspective, combining a recurrent VAE prior with a Gaussian likelihood model. Unsupervised dereverberation with a non-generative prior has also been investigated in the multi-channel scenario [39].
This paper expands our prior work [40], where we designed a blind unsupervised dereverberation algorithm, extending [26] to the blind scenario. The resulting approach, called BUDDy, uses a model-based parametric subband filter with an exponential decay to approximate the RIR. BUDDy performs joint estimation of the RIR and the anechoic speech, leveraging the model-based parameterization as an acoustic prior and the diffusion model as a speech prior. We have shown previously [40] that BUDDy can successfully remove reverberation, and that it is robust to changes in acoustic conditions because of the lack of supervision during training. Therefore, BUDDy closes the performance gap between matched and mismatched acoustic conditions in comparison to diffusion-based supervised approaches [24, 25].
In this paper, we extend the experimental framework of our previous publication [40] with the following contributions:
  • Section IV-A extends the evaluation of BUDDy for speech dereverberation beyond instrumental metrics, including a subjective listening test and a set of ablation studies. Notably, we perform experiments on speech recordings made in real environments, rather than using the convolution model, a scenario in which BUDDy performs especially well.
  • Section IV-B presents new experiments on applying BUDDy to singing voice dereverberation at a sampling rate of 44.1 kHz, which is higher than the 16-kHz sampling rate used in our speech experiments [40]. The results, which also include a subjective listening test, indicate that our method significantly outperforms existing unsupervised state-of-the-art approaches and performs comparably to supervised baselines.
  • In Section IV-C, we investigate the robustness of informed dereverberation approaches in partially blind scenarios, in comparison to BUDDy. We highlight the limitations of these approaches when the RIR is perturbed with Gaussian noise or estimated blindly using a state-of-the-art RIR estimator [41].
  • Finally, Section IV-D assesses BUDDy’s performance in RIR estimation against a state-of-the-art supervised estimator [41]. We use frequency-wise acoustic descriptors to evaluate the accuracy of BUDDy on reverberation time and clarity.
We organize the paper as follows. In Section II, we introduce diffusion-based generative models and posterior sampling methods for informed dereverberation using diffusion priors as proposed in previous work [26]. Then in Section III, we introduce our blind unsupervised dereverberation method BUDDy [40], which extends the posterior sampling method presented in the previous section to the blind scenario where the RIR is not available. The experiments and results mentioned above are presented in Section IV. Section V concludes the paper.
II Informed Diffusion-Based Dereverberation
This section introduces diffusion models, a class of generative models that form the foundation of the proposed method. It also explores their application in solving inverse problems, specifically highlighting their use in informed dereverberation.
Throughout this paper, we use the following notations: considering dereverberation under the prism of inverse problem solving, we wish to retrieve the anechoic time-domain utterance $\mathbf{x}_{0}\in\mathbb{R}^{L}$, where $L$ is the length of the utterance, given the reverberant measurement $\mathbf{y}$. As in most dereverberation studies, reverberation is modeled as a convolution between anechoic speech with a RIR $\mathbf{h}\in\mathbb{R}^{L_{\mathbf{h}}}$, such that $\mathbf{y}=\mathbf{h}\ast\mathbf{x}_{0}$, where $\ast$ is the discrete convolution operator in the time domain, resulting in $\mathbf{y}\in\mathbb{R}^{L+L_{\mathbf{h}}-1}$.
II-A Diffusion-Based Generative Models
Diffusion models [22, 42] have achieved remarkable success across various domains, including speech [43]. They break down the problem of generating high-dimensional complex data into a series of easier denoising tasks. Training a diffusion model first requires defining a forward process, which gradually adds noise to data points, turning the target data distribution into a tractable Gaussian distribution.
The forward process is the solution of the following stochastic differential equation (SDE):
$\mathrm{d}\mathbf{x}_{\tau}=\mathbf{f}(\mathbf{x}_{\tau},\tau)\mathrm{d}\tau+g(\tau)\mathrm{d}\mathbf{w}_{\tau}\,,$ (1)
where the diffusion time $\tau$ indexes the stochastic process $\mathbf{x}_{\tau}$ from $\tau=T_{\mathrm{min}}$ to $\tau=T\gg T_{\mathrm{min}}$. The minimal process time $T_{\mathrm{min}}$ is chosen stricly positive to avoid irregularities around 00. The Wiener process $\mathbf{w}_{\tau}$ injects noise with independent and normally distributed increments, that is, $\mathbf{w}_{\tau+\mathop{}\!\mathrm{d}{\tau}}-\mathbf{w}_{\tau}\sim\mathcal{N}(\mathbf{0},\mathop{}\!\mathrm{d}{\tau}\,\mathbf{I})$ where $\mathbf{I}\in\mathbb{R}^{L\times L}$ is the identity matrix [44]. The diffusion state $\mathbf{x}_{\tau}\in\mathbb{R}^{L}$ starts at a clean speech data point $\mathbf{x}_{0}\in\mathbb{R}^{L}\sim p_{\text{data}}$ and ends at the final state $\mathbf{x}_{T}\in\mathbb{R}^{L}$ which contains mostly Gaussian noise. We adopt the parameterization proposed by Karras et al. [45], which defines the drift and diffusion parameters as $f(\mathbf{\mathbf{x}_{\tau}},\tau)=0$ and $g(\tau)=\sqrt{2\tau}$, respectively. This results in a noise schedule $\sigma(\tau)=\tau$ which determines the so-called transition kernel i.e. the marginal density of the forward process [45]
$q_{\tau}(\mathbf{x}_{\tau}|\mathbf{x}_{0})=\mathcal{N}(\mathbf{x}_{0},\sigma^{2}(\tau)\mathbf{I})\,.$ (2)
Conversely, data generation is accomplished by reversing the forward corruption process. First, an initial sample is drawn from a Gaussian distribution, and then the model iteratively removes noise until a clean sample from the target distribution emerges. The reverse process, can be characterized by the probability flow ordinary differential equation (ODE), which has the same marginal distributions as the reverse SDE canonically associated to the forward SDE (1) [46]
$\mathrm{d}\mathbf{x}_{\tau}=\left[\mathbf{f}(\mathbf{x}_{\tau},\tau)-\frac{g^{2}(\tau)}{2}\nabla_{\mathbf{x}_{\tau}}\log p(\mathbf{x}_{\tau})\right]\mathrm{d}\tau\,,$ (3)
where diffusion time $\tau$ flows in reverse from $\tau=T$ to $\tau=T_{\mathrm{min}}$. The diffusion state $\mathbf{x}_{\tau}\in\mathbb{R}^{L}$ starts from the initial state $\mathbf{x}_{T}\in\mathbb{R}^{L}$ and ends at $\mathbf{x}_{0}\in\mathbb{R}^{L}\sim p_{\text{data}}$. The score function $\nabla_{\mathbf{x}_{\tau}}\log p(\mathbf{x}_{\tau})$ indicates the direction towards regions of higher probability under the model’s distribution. In practice, it is intractable and we need to estimate it with a score model $\mathbf{s}_{\theta}(\mathbf{x}_{\tau},\tau)$ parameterized with a deep neural network (DNN). Vincent et al. have shown that the score model $\mathbf{s}_{\theta}(\mathbf{x}_{\tau},\tau)$ can be optimized using denoising score matching [47], i.e. matching the score of the Gaussian transition kernel $q_{\tau}(\mathbf{x}_{\tau}|\mathbf{x}_{0})$ instead of the score of the unknown probability $p(\mathbf{x}_{\tau})$. The score of the transition kernel $q_{\tau}(\mathbf{x}_{\tau}|\mathbf{x}_{0})$ can be obtained from (2) as
$\nabla_{\mathbf{x}_{\tau}}\log q_{\tau}(\mathbf{x}_{\tau}|\mathbf{x}_{0})=-\frac{\mathbf{x}_{\tau}-\mathbf{x}_{0}}{\sigma^{2}(\tau)}\,.$ (4)
The score model $\mathbf{s}_{\theta}$ is therefore trained using the denoising score-matching objective [47]
[…]
x_τ∼q_τ( x_τx_0)
[…]
x_τ∼q_τ( x_τx_0)
[…]
where first a diffusion index $\tau$ is randomly sampled between extremal times $T_{\mathrm{min}}$ and $T_{\mathrm{max}}>T$, a data point $\mathbf{x}_{0}$ is sampled in the training set, and the corresponding diffusion state $\mathbf{x}_{\tau}$ is obtained from the transition kernel in (2). In practice, we use the same pre-conditioning for $\mathbf{s}_{\theta}(\mathbf{x}_{\tau},\tau)$ and same loss weighting $\lambda(\cdot)$ as in Karras et al. (see [45] for details).
II-B Diffusion Posterior Sampling for Dereverberation
We discuss in this section how diffusion priors can be adapted in order to solve inverse problems. While some traditional methods derive maximum a posteriori estimators for blind dereverberation [9, 10, 11], we exploit the generative nature of diffusion models to solve this inverse problem using posterior sampling. Assuming that the RIR $\mathbf{h}$ is known, we attempt to sample from the posterior distribution of the anechoic speech given the measurement and the RIR $p(\mathbf{x}_{0}|\mathbf{y})\propto p(\mathbf{y}|\mathbf{x}_{0})p(\mathbf{x}_{0})$. Although we do not have an explicit prior $p(\mathbf{x}_{0})$ like in e.g. VAE frameworks, we leverage the implicit prior $p_{\theta^{*}}(\mathbf{x}_{0})$ given by the pretrained diffusion model $\mathbf{s}_{\theta^{*}}(\mathbf{x}_{\tau},\tau)$ where $\theta^{*}$ represents the (fixed) parameters optimizing the training objective (5). Sampling is then achieved by solving the probability flow ODE (3), replacing the unconditional score function by the posterior score $\nabla_{\mathbf{x}_{\tau}}\log p(\mathbf{x}_{\tau}|\mathbf{y})$ [23] obtained through Bayes’ rule
$\mathrm{d}\mathbf{x}_{\tau}$ $=\left[\mathbf{f}(\mathbf{x}_{\tau},\tau)-\frac{1}{2}g^{2}(\tau)\nabla_{\mathbf{x}_{\tau}}\log p(\mathbf{x}_{\tau}|\mathbf{y})\right]\mathrm{d}\tau\,,$
$\approx\left[\mathbf{f}\left(\mathbf{x}_{\tau},\tau\right)-\frac{1}{2}g^{2}(\tau)\left(\mathbf{s}_{\theta^{*}}(\mathbf{x}_{\tau},\tau)+\nabla_{\mathbf{x}_{\tau}}\log p(\mathbf{y}|\mathbf{x}_{\tau})\right)\right]\mathrm{d}\tau\,.$ (6)
The likelihood score $\nabla_{\mathbf{x}_{\tau}}\log p(\mathbf{y}|\mathbf{x}_{\tau})$ is in general intractable for $\tau>0$. Following [27], we employ a plug-in estimate of $\mathbf{x}_{0}$ denoted as $\hat{\mathbf{x}}_{0}^{\tau}$ which we derive using Tweedie’s formula, i.e. one-step denoising of $\mathbf{x}_{\tau}$ using the diffusion model
$\hat{\mathbf{x}}_{0}^{\tau}\overset{\Delta}{=}\mathbb{E}[\mathbf{x}_{0}|\mathbf{x}_{\tau}]\approx\mathbf{x}_{\tau}+\sigma^{2}(\tau)\mathbf{s}_{\theta^{*}}(\mathbf{x}_{\tau},\tau)\,.$ (7)
We assume that this estimate is a sufficient statistic for $\mathbf{x}_{\tau}$, which results in a first assumption $p(\mathbf{y}|\mathbf{x}_{\tau})\approx p(\mathbf{y}|\hat{\mathbf{x}}_{0}^{\tau})$.
In order to approximate $p(\mathbf{y}|\hat{\mathbf{x}}_{0}^{\tau})$, previous work [26] models the error between $\mathbf{y}$ and its estimation to follow a zero-mean Gaussian distribution in the time domain. The corresponding expression for the likelihood score $\nabla_{\mathbf{x}_{\tau}}\log p(\mathbf{y}|\mathbf{x}_{\tau})$ is then a simple weighted $L^{2}$-distance between $\mathbf{y}$ and $\mathbf{h}\ast\hat{\mathbf{x}}_{0}^{\tau}$. However far better dereverberation performance and speech quality can be achieved by substituting the obtained distance with a $L^{2}$-distance between compressed short-time Fourier transform (STFT) representations instead. This is analogous to modeling the likelihood as
$p(\mathbf{y}|\mathbf{x}_{0})\propto\mathrm{exp}\big{(}-\zeta(\tau)\,\mathcal{C}(\mathbf{y},\mathbf{h}\ast\hat{\mathbf{x}}_{0}^{\tau})\,\big{)}\,,$ (8)
where the scalar $\zeta(\tau)$ controls the influence of the likelihood score term during sampling, and $\mathcal{C}(\cdot,\cdot)$ is the cost function
$\mathcal{C}(\mathbf{u},\mathbf{v})=\frac{1}{M}\sum_{m=1}^{M}\sum_{k=1}^{K}\rVert S_{\mathrm{comp}}(\mathbf{u})_{m,k}-S_{\mathrm{comp}}(\mathbf{v})_{m,k}\lVert_{2}^{2}\,.$ (9)
There, $S_{\mathrm{comp}}(\mathbf{u})\in\mathbb{C}^{M\times K}$ denotes the magnitude-compressed STFT of $\mathbf{u}$, comprising $M$ time frames and $K$ frequency bins
$S_{\mathrm{comp}}(\mathbf{u})=|\text{STFT}(\mathbf{u})|^{2/3}\exp{j\angle\text{STFT}(\mathbf{u})}\,.$ ) . (10)
We apply this compression to boost low-energy components as typically observed in high frequencies of speech signals or in late reverberation tails, and account for the heavy-tailedness of speech distributions [4]. Such a strategy is also employed in [48] for data representation. In our case, we only use this non-linear transformation in the cost function (9), whereas the diffusion process itself uses the original time-domain representation.
The parameter $\zeta(\tau)$ balances a trade-off between adherence to the prior data distribution and fidelity to the observed data. We empirically resort to the same parameterization of $\zeta(\tau)$ as in [49, 32]:
$\zeta(\tau)=\frac{\sqrt{L}\,\tilde{\zeta}}{\sigma(\tau)\norm{\nabla_{\mathbf{x}_{\tau}}\mathcal{C}\left(\mathbf{y},\mathbf{h}\ast\hat{\mathbf{x}}_{0}(\mathbf{x}_{\tau})\right)}_{2}}\,,$ (11)
where $\tilde{\zeta}$ is a fixed coefficient.
The resulting informed dereverberation algorithm is a slight variation of our previous work [26]. In the following we refer to this approach as InfDerevDPS.
III Blind Diffusion-Based Dereverberation
Algorithm 1 Reverberation Operator $\mathcal{A}_{\psi}(\cdot)$
function $\mathcal{A}_{\psi}$($\hat{\mathbf{x}}_{0}$)
     $\{\mathbf{\Phi},(w_{b},\alpha_{b})_{b=1,\dots,B}\}\leftarrow\psi$triangleright Parameter set
     $\mathbf{A}_{n,b}^{\prime}\leftarrow w_{b}\cdot e^{-\alpha_{b}n}$triangleright Exponential decay model
     $\mathbf{A}\leftarrow\exp{\left(\mathrm{lerp}(\log{\mathbf{A}^{\prime}})\right)}$ ) $\triangleright$ Frequency interpolation
     $\mathbf{H}\leftarrow\mathbf{A}\cdot e^{j\mathbf{\Phi}}$
     $\overline{\mathbf{H}}\leftarrow\mathrm{STFT}(\delta_{\text{d}}\oplus\mathcal{P}_{\text{min}}(\mathrm{iSTFT}(\mathbf{H})))$triangleright Projection step
     $\hat{\mathbf{X}}\leftarrow\mathrm{STFT}(\hat{\mathbf{x}}_{0})$
     $\hat{\mathbf{Y}}_{m,k}\leftarrow\sum_{n=0}^{N_{h}}\overline{\mathbf{H}}_{n,k}\hat{\mathbf{X}}_{m-n,k}$triangleright Subband convolution
     return $\mathrm{iSTFT}(\hat{\mathbf{Y}})$
Algorithm 2 Inference algorithm
Reverberant speech $\mathbf{y}$
$\mathbf{x}_{\mathrm{init}}\leftarrow\mathrm{{\color[rgb]{0.52734375,0.30078125,0.75}\definecolor[named]{pgfstrokecolor}{rgb}{0.52734375,0.30078125,0.75}WPE}}(\mathbf{y})$
Sample $\mathbf{x}_{N}\sim\mathcal{N}({\mathbf{x}_{\mathrm{init}}},\sigma_{N}^{2}\mathbf{I})$triangleright Warm initialization
Initialize $\psi_{N}$triangleright Initialize the RIR parameters
for $n\leftarrow N,\dots,1$ do $\triangleright$ Discrete step backwards
   Phase 1 – E-step
     $\mathbf{s}_{n}\leftarrow s_{\theta}(\mathbf{x}_{n},\tau_{n})$triangleright Evaluate score model
     $\hat{\mathbf{x}}_{0}\leftarrow\mathbf{x}_{n}-\sigma_{n}^{2}\mathbf{s}_{n}$triangleright Get one-step denoising estimate
     $\hat{\mathbf{x}}_{0}\leftarrow\mathrm{Rescale}(\hat{\mathbf{x}}_{0})$triangleright Constraint RMS power
     $\mathbf{g}_{n}\leftarrow-\zeta(\tau_{n})\nabla_{\mathbf{x}_{n}}\mathcal{C}(\mathbf{y},\mathcal{A}_{\psi_{n}}(\hat{\mathbf{x}}_{0}))$triangleright LH score approx.
     $\mathbf{x}_{{n-1}}\leftarrow\mathbf{x}_{n}-\sigma_{n}(\sigma_{n-1}-\sigma_{n})(\mathbf{s}_{n}+\mathbf{g}_{n})$triangleright Update step
   Phase 2 – M-step
     $\psi_{n-1}^{0}\leftarrow\psi_{n}$triangleright Use RIR parameters from last step
     for $j\leftarrow 0,\dots,N_{\text{its.}}$ do $\triangleright$ RIR optimization
         $\mathcal{J}_{\text{RIR}}(\psi_{n-1}^{j})\leftarrow\mathcal{C}(\mathbf{y},\mathcal{A}_{\psi_{n-1}^{j}}(\hat{\mathbf{x}}_{0}))+\mathcal{R}(\psi_{n-1}^{j})$
         $\psi_{n-1}^{j+1}\leftarrow\psi_{n-1}^{j}-\mathrm{Adam}(\mathcal{J}_{\text{RIR}}(\psi_{n-1}^{j}))$triangleright Opti. step
         $\psi_{n-1}^{j+1}\leftarrow\mathrm{clamp}(\psi_{n-1}^{j+1})$triangleright Constrain Parameters
     $\psi_{n-1}\leftarrow\psi_{n-1}^{M}$
return $\mathbf{x}_{0}$triangleright Reconstructed audio signal
$\mathbf{x}_{N}$$\psi_{N}$
$\mathbf{x}_{n}$$\psi_{n}$
[…]
Figure 1: BUDDy: joint optimization alternating between RIR estimation and posterior sampling for speech reconstruction [40].
This section elaborates on the proposed method BUDDy, which extends the informed method presented in prior work [26] to the blind scenario, where the impulse response $\mathbf{h}$ is unknown. In Section III-A, we define a reverberation operator $\mathcal{A}_{\psi}(\cdot)$, which comprises a structured parametric model of the RIR, with parameters $\psi$. Section III-B then details the posterior sampling scheme used to obtain a speech utterance from the estimated posterior distribution. Finally, Section III-C describes the optimization procedure for the reverberation model parameters $\psi$ using an expectation-maximization (EM) formulation. The complete inference procedure is summarized in Algorithm 2, and an overview of the processing pipeline is given in Fig. 1.
III-A Reverberation Operator
III-A1 Subband Filtering
In contrast to [26] where the original time-domain convolution model is adopted, here we model reverberation using a subband filtering approximation in the STFT domain [50, 51]. This approach enables us to incorporate prior knowledge about the characteristics of reverberation through a structured model of RIR magnitudes, characterized by exponential decays in each subband. By modeling reverberation in this way, we facilitate optimization and reduce the complexity of the RIR parameter search space. Let $\mathbf{H}:=\mathrm{STFT}(\mathbf{h})\in\mathbb{C}^{N_{\mathbf{h}}\times K}$ represent the STFT of a RIR $\mathbf{h}$ with $N_{\mathbf{h}}$ time frames and $K$ frequency bins. Similarly, let $\mathbf{X}\in\mathbb{C}^{M\times K}$, and $\mathbf{Y}$, denote the STFTs of anechoic $\mathbf{x}_{0}$ and reverberant $\mathbf{y}$ speech signals, repectively. The subband convolution operation applies independent convolutions along the time dimension of each frequency band:
$\mathbf{Y}_{m,k}=\sum_{n=0}^{N_{h}}\mathbf{H}_{n,k}\mathbf{X}_{m-n,k}\,.$ (12)
The resulting reverberant signal $\mathbf{Y}\in\mathbb{C}^{(M+N_{\mathbf{h}}-1)\times K}$ can be transformed to time domain by applying the inverse STFT. The subband filtering model only approximates the time-domain convolution, as it does not account for the spectral leakage between frequency bands. However, it is empirically found to be a valid assumption in many scenarios involving reverberation [51, 7, 52]. Adding 50% zero-padding to the end of the frames before computing the STFT is important to avoid cyclic convolution artifacts when transforming the resulting signal back to the time domain.
III-A2 RIR Prior
In the blind scenario, estimating $\mathbf{H}$ is an ill-posed problem when the anechoic speech is unknown. Therefore, we need to constrain the space of possible solutions by imposing a prior on $\mathbf{H}$. We propose a structured, differentiable prior on $\mathbf{H}$, whose parameters $\psi$ can be estimated with gradient-based optimizers like Adam [53]. We denote the complete forward reverberation operator, including forward and inverse STFT operations, as $\mathcal{A}_{\psi}(\cdot):\mathbb{R}^{L}\rightarrow\mathbb{R}^{L+L_{\mathbf{h}}-1}$. The whole processing pipeline is summarized in Algorithm 1 with each component detailed below.
We denote as $\mathbf{A}\in\mathbb{R}^{N_{\mathbf{h}}\times K}$ and $\mathbf{\Phi}\in\mathbb{R}^{N_{\mathbf{h}}\times K}$ the RIR magnitudes and phases, respectively. Following [11], we adopt an exponential decay model with learnable parameters controlling the decay time. Since room materials exhibit frequency-dependent absorption behavior, we parameterize the magnitude matrix $\mathbf{A}$ as a multi-band exponential decay model defined in $B<K$ frequency bands. Let $\mathbf{A}^{\prime}\in\mathbb{R}^{N_{\mathbf{h}}\times B}$ be the subsampled version of $\mathbf{A}$ in the $B$ selected frequency bands. Each frequency band $b$ is characterized by its weight $w_{b}$ and exponential decay rate $\alpha_{b}$, such that the corresponding subband magnitude filter is derived as
$\mathbf{A}_{n,b}^{\prime}=w_{b}\cdot e^{-\alpha_{b}n}\,.$ (13)
Note that our parameterization can be extended to model coupled spaces by employing several decay parameters per band and summing their respective contributions [54]. Once the parameters are estimated, we reconstruct the $K$-bands magnitudes $\mathbf{A}$ by interpolating the subsampled matrix $\mathbf{A}^{\prime}$ as $\mathbf{A}=\exp(\mathrm{lerp}(\log(\mathbf{A}^{\prime})))$, where $\mathrm{lerp}$ represents linear interpolation on the frequency scale. For this purpose, we employ the $\mathrm{torchcde}$ library, which facilitates efficient and differentiable interpolation [55]. After interpolation of the magnitude matrix, we then obtain the time-frequency RIR $\mathbf{H}$ by multiplying the magnitude matrix $\mathbf{A}$ with the complex phase exponentials:
$\mathbf{H}=\mathbf{A}\odot e^{j\mathbf{\Phi}},$ (14)
where $j$ is the imaginary number and $\odot$ represents element-wise multiplication. Given the general lack of phase structure, we optimize each phase factor in $\mathbf{\Phi}$ independently. The RIR model $\psi=\{\mathbf{\Phi},(w_{b},\alpha_{b})_{b=1,\dots,B}\}$ ultimately contains $2\times B+N_{\mathbf{h}}\times K$ optimizable parameters.
III-A3 Projections
We extend our forward reverberation operator with a series of projections to increase the likelihood of generating plausible RIRs. Thus, the time-frequency RIR $\mathbf{H}$ is further processed as
$\overline{\mathbf{H}}=\mathrm{STFT}\left(\delta\oplus\mathcal{P}_{\text{min}}(\mathrm{iSTFT}(\mathbf{H}))\right)\,.$ (15)
This primarily ensures STFT consistency of $\overline{\mathbf{H}}$, exploiting the redundancy of the STFT representation and imposing inter-frame correlations between the RIR phases $\boldsymbol{\Phi}$. We then enforce that the time-domain RIR estimate $\mathbf{h}$ has minimum-phase lag, using the Hilbert transform-based method in [56]. This is indicated by the operator $\mathcal{P}_{\text{min}}$ and guarantees stability of the inverse RIR filter [2]. We refer the reader to Appendix -A for further details. Finally, the operation $\delta\oplus(\cdot)$ replaces the first sample of the time-domain RIR with a unit impulse. This has the effect of injecting knowledge of the direct path in $\overline{\mathbf{H}}$, and further requires us to correct the magnitude matrix $\mathbf{A}$ to account for this operation. It is important to note that these steps are integral to the reverberation operator $\mathcal{A}_{\psi}(\cdot)$, which maps the parameters $\psi$ to the convolved signal $\mathcal{A}_{\psi}(\hat{\mathbf{x}}_{0})$, as outlined in Algorithm 1. Since all operations are differentiable, we compute gradients with respect to $\psi$ by backpropagating through all operations. We propose a detailed ablation study of these projection and correction steps in Section IV-A6.
III-B Posterior Speech Sampling
For sampling a speech utterance from the posterior distribution $p_{\psi}(\mathbf{x}_{0}|\mathbf{y})$, we adapt the posterior sampling algorithm of Section II-B to the blind setting. As in Section II-B, the pre-trained score model $\mathbf{s}_{\theta^{\ast}}(\mathbf{x}_{\tau},\tau)$ is used to provide an implicit prior $p_{\theta^{\ast}}(\mathbf{x}_{0})$ on anechoic speech with fixed parameters $\theta^{\ast}$. The likelihood $p_{\psi}(\mathbf{y}|\mathbf{x}_{0})$ is approximated as
$p_{\psi}(\mathbf{y}|\mathbf{x}_{0})\propto\mathrm{exp}\left(-\zeta(\tau)\mathcal{C}(\mathbf{y},\mathcal{A}_{\psi}(\hat{\mathbf{x}}_{0}^{\tau})\right)\,,$ (16)
which is analogous to (8), but with the proposed reverberation operator $\mathcal{A}_{\psi}(\cdot)$ applied in place of a convolution with the oracle RIR, since the latter is unavailable in our blind scenario. The sampling procedure then follows the ODE (II-B) where the likelihood term $p_{\psi}(\mathbf{y}|\mathbf{x}_{0})$ defined in (16) is used instead of the RIR-informed likelihood $p(\mathbf{y}|\mathbf{x}_{0})$ from (8).
In order to guide and accelerate reverse diffusion, it is beneficial to use warm initialization, i.e., let the reverse diffusion process start from a speech sample $\mathbf{x}_{T}\sim\mathcal{N}(\mathbf{x}_{\text{init}},\sigma^{2}(T)\mathbf{I})$, where $\mathbf{x}_{\text{init}}$ contains relevant information about the clean signal we wish to estimate. Similar to [34], we obtain $\mathbf{x}_{\text{init}}$ through WPE [7], a blind dereverberation algorithm based on variance-normalized delayed linear prediction. WPE performs mild dereverberation, which allows us to get closer to the clean speech, while not introducing much distortion to the signal. As WPE is blind and unsupervised, our method remains fully blind and unsupervised as well.
III-C Reverberation Model Parameter Optimization
The posterior speech sampling approach described in Section III-B relies on approximating the likelihood $p_{\psi}(\mathbf{y}|\mathbf{x}_{0})$, where the main challenge is the dependence on the unknown parameters $\psi$. To estimate these parameters, we employ an EM formulation, drawing inspiration from [31, 29]. While the underlying algorithm follows a similar approach to prior works [57, 40], the formalism has been refined to offer a more precise explanation of the method. The optimization procedure maximizes the expected log-likelihood $\log p_{\psi}(\mathbf{y}|\mathbf{x}_{0})$ under the posterior distribution $p_{\psi}(\mathbf{x}_{0}|\mathbf{y})$
[…]
This objective is unfortunately intractable, because the search quantity $\psi$ appears as a parameter of the distribution used to evaluate the expectation. Therefore, we resort to an EM formulation alternating between an E-step and an M-step. During the E-step, the expectation is evaluated by drawing samples from the approximate posterior $p_{\psi}(\mathbf{x}_{0}|\mathbf{y})$, given fixed RIR parameters $\psi$. The M-step then optimizes the RIR parameters $\psi$ based on the clean speech estimates, maximizing the expected log-likelihood. Both steps are detailed below.
III-C1 E-Step
Given RIR parameters $\psi$, we wish to approximate the posterior speech distribution $p_{\psi}(\mathbf{x}_{0}|\mathbf{y})$ in order to evaluate the expectation in (17). Drawing samples through our posterior sampling procedure explained above in Section III-B would require running the entire diffusion process from $T$ to $T_{\mathrm{min}}$ at each EM iteration, as in [31]. Instead, we follow [29] and evaluate the expectation using the one-step denoised estimate $\hat{\mathbf{x}}_{0}^{\tau}$ through (7) that is obtained at each reverse diffusion step. Formally, the log-likelihood expectation is approximated as
$\mathbb{E}_{p_{\psi}(\mathbf{x}_{0}|\mathbf{y})}\log p_{\psi}(\mathbf{y}|\mathbf{x}_{0})\approx\mathbb{E}_{p(\mathbf{x}_{0}|\mathbf{x}_{\tau})}\log p_{\psi}(\mathbf{y}|\mathbf{x}_{0})\,,$ (18)
and the denoised posterior $p(\mathbf{x}_{0}|\mathbf{x}_{\tau})$ is modelled as a Dirac distribution located at the posterior mean $\hat{\mathbf{x}}_{0}^{\tau}$ [27]
$p(\mathbf{x}_{0}|\mathbf{x}_{\tau})\approx\delta(\hat{\mathbf{x}}_{0}^{\tau})\,.$ (19)
This results in the following one-sample Monte Carlo estimate of the expectation (17)
$\mathbb{E}_{p_{\psi}(\mathbf{x}_{0}|\mathbf{y})}\log p_{\psi}(\mathbf{y}|\mathbf{x}_{0})\approx\log p_{\psi}(\mathbf{y}|\hat{\mathbf{x}}_{0}^{\tau})\,.$ (20)
This effectively integrates the M-step into the reverse diffusion, significantly accelerating inference. Note that in (18), the dependency of $p_{\psi}(\mathbf{x}_{0}|\mathbf{y})$ to $\mathbf{y}$ is no longer explicit. However, this dependency still exists, as we notice that the Tweedie estimate $\hat{\mathbf{x}}_{0}^{\tau}$ is obtained from the current diffusion state $\mathbf{x}_{\tau}$, which has been itself sampled conditionally to $\mathbf{y}$ in (II-B).
III-C2 M-step
Once the expectation in (17) has been evaluated during the E-step, we can proceed to the M-step that will maximize the resulting objective. We add to the expected log-likelihood a regularization term facilitating the RIR parameter search during optimization
$\mathcal{R}(\psi)=\frac{1}{N_{\mathbf{h}}}\rVert S_{\mathrm{comp}}(\mathbf{h}_{\psi})-S_{\text{comp}}(\mathrm{sg}\left[\mathbf{h}_{\psi}\right]+\nu(\tau)\mathbf{v})\ \lVert_{2}^{2},$ (21)
where $\mathbf{h}_{\psi}=\mathcal{A}_{\psi}(\delta)$ is the current time-domain RIR estimate, $\mathbf{v}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ is a vector of white Gaussian noise, and $S_{\mathrm{comp}}$ is the magnitude compressed spectrogram, as defined in (10), with $N_{\mathbf{h}}$ time frames. In the right-hand term, the “stopgrad” operator $\mathrm{sg}[\cdot]$ detaches the gradients of $\mathbf{h}_{\psi}$ from the optimization graph and Gaussian noise scaled by $\nu(\tau)$ is added. We show in Appendix -B that this effectively injects multiplicative noise with standard deviation $\nu(\tau)$ in the RIR parameter gradients, taking inspiration from the “regularization by denoising” strategy used, e.g. in [58] for guiding data reconstruction. In our case, this regularization smoothes the RIR parameter optimization landscape by injecting small amounts of stochasticity during the optimization.
The M-step finally consists in maximizing the resulting objective:
[…]
Since all operations, including our reverberation operator $\mathcal{A}_{\psi}(\cdot)$, are differentiable with respect to $\psi$, the M-step can be achieved with gradient-based optimizers, e.g. Adam [53].
During optimization, we further rescale the denoised speech estimate $\hat{\mathbf{x}}_{0}^{\tau}$ so that its root-mean-square power (RMS) matches the average RMS power of clean speech computed on the training set. Using this additional constraint helps lift the indeterminacy when jointly optimizing the speech $\mathbf{x}_{0}$ and RIR parameters $\psi$. This step is included in our ablation study in Section IV-A6. We also found it beneficial to constrain $w_{b}$ and $\alpha_{b}$ within a limited range to stabilize the optimization, specially at early stages. This is achieved by clamping the parameters to predefined minimum and maximum values after every optimization iteration, as specified in Appendix -C1.
IV Experiments and Results
In this section, we provide a comprehensive evaluation of BUDDy across various datasets and experimental setups. We detail the methodologies and baselines employed and present the results of our experiments.
IV-A Speech Dereverberation
We present dereverberation results on 16 kHz speech data, building upon the experiments conducted in prior work [40].
TABLE I: Speech dereverberation results on reverberant VCTK datasets. We indicate for each method in the table whether it is supervised or not. Boldface numbers indicate best performance for supervised and unsupervised methods separately.
VCTK-RealReverb (Matched) VCTK-SimulatedReverb (Mismatched)
Method Unsup. DNS-MOS PESQ ESTOI SI-SDR DNS-MOS PESQ ESTOI SI-SDR
Reverberant - 3.14 $\pm$ 0.52 1.61 $\pm$ 0.37 0.50 $\pm$ 0.14 -12.3 $\pm$ 6.9 3.05 $\pm$ 0.47 1.57 $\pm$ 0.29 0.47 $\pm$ 0.11 -12.5 $\pm$ 8.6
PSE 3.75 $\pm$ 0.38 2.85 $\pm$ 0.55 0.80 $\pm$ 0.10 8.5 $\pm$ 6.5 3.61 $\pm$ 0.39 2.08 $\pm$ 0.47 0.64 $\pm$ 0.09 -8.4 $\pm$ 8.6
SGMSE+M [24, 59] 3.88 $\pm$ 0.32 2.99 $\pm$ 0.48 0.78 $\pm$ 0.09 0.2 $\pm$ 9.3 3.74 $\pm$ 0.34 2.48 $\pm$ 0.47 0.69 $\pm$ 0.09 -8.4 $\pm$ 8.8
StoRM [25] 3.90 $\pm$ 0.33 3.33 $\pm$ 0.48 0.82 $\pm$ 0.10 9.5 $\pm$ 6.5 3.83 $\pm$ 0.32 2.51 $\pm$ 0.53 0.67 $\pm$ 0.09 -8.8 $\pm$ 10.2
Yohena and Yatabe [8] 2.99 $\pm$ 0.56 1.80 $\pm$ 0.33 0.55 $\pm$ 0.12 -11.4 $\pm$ 7.7 2.94 $\pm$ 0.44 1.71 $\pm$ 0.29 0.51 $\pm$ 0.10 -11.4 $\pm$ 8.6
WPE [60] 3.24 $\pm$ 0.54 1.81 $\pm$ 0.42 0.57 $\pm$ 0.14 -11.5 $\pm$ 8.2 3.10 $\pm$ 0.48 1.74 $\pm$ 0.37 0.54 $\pm$ 0.12 -11.4 $\pm$ 8.8
Saito et al. [34] 3.22 $\pm$ 0.56 1.68 $\pm$ 0.40 0.51 $\pm$ 0.13 -11.7 $\pm$ 9.2 3.12 $\pm$ 0.52 1.70 $\pm$ 0.33 0.52 $\pm$ 0.10 -11.7 $\pm$ 8.5
GibbsDDRM [35] 3.33 $\pm$ 0.53 1.70 $\pm$ 0.37 0.51 $\pm$ 0.13 -11.9 $\pm$ 8.5 3.30 $\pm$ 0.52 1.75 $\pm$ 0.36 0.52 $\pm$ 0.11 -11.8 $\pm$ 8.9
RVAE-EM [38] 3.05 $\pm$ 0.53 1.83 $\pm$ 0.32 0.54 $\pm$ 0.11 -12.2 $\pm$ 8.1 3.00 $\pm$ 0.45 1.76 $\pm$ 0.30 0.52 $\pm$ 0.10 -11.8 $\pm$ 8.5
BUDDy (ours) 3.76 $\pm$ 0.41 2.30 $\pm$ 0.53 0.66 $\pm$ 0.12 -7.8 $\pm$ 8.40 3.74 $\pm$ 0.38 2.24 $\pm$ 0.54 0.65 $\pm$ 0.12 -8.4 $\pm$ 9.9
IV-A1 Data
We use VCTK [61] as clean speech, selecting 103 speakers for training, two for validation, and two for testing. The total dataset represents 44 h of audio, which we down-sample to 16 kHz for our experiments. We curate RIRs from various public datasets [62, 63, 64, 65, 66, 67, 68, 69, 70]. In total we approximately obtain 10k RIRs, and split them between training, validation and testing using ratios 0.9/0.05/0.05. We use the three following test benchmarks for our experiments:
[…]
  • VCTK-SimulatedReverb: This test set uses the same test speech utterances as VCTK-RealReverb, matching the training speech corpus. However, the reverberant conditions are obtained by simulating RIRs with pyroomacoustics[71]. For ease of comparison, we choose simulation parameters such that the distributions of reverberation times and direct-to-reverberation ratios of the simulated mismatched dataset approximately match those of the matched dataset using real RIRs.
[…]
IV-A2 Baselines
We compare our method BUDDy to several blind supervised baselines such as the predictive approach in [59], which will denote as PSE in the following (for predictive speech enhancement), and diffusion-based SGMSE+ [24] and StoRM [25]. The STFT-based diffusion model in SGMSE+ and StoRM uses supervision in both the network conditioning and the diffusion trajectory parameterization; PSE uses a classical $L^{2}$-distance between the clean target and its estimate and has virtually the same architecture as SGMSE+. These methods require coupled reverberant/anechoic speech, which we generate using our curated RIR and anechoic speech datasets. The reverberant speech is obtained by first aligning the direct path of the RIR to its first sample, then convolving the anechoic speech from VCTK with the resulting RIR, and finally normalizing it to reach the same loudness [74] as the anechoic speech.
We also include blind unsupervised approaches leveraging traditional methods such as WPE [7] and Yohena and Yatabe [8], as well as generative models Saito et al. [34], GibbsDDRM [35] and RVAE-EM [38]. Please see Appendix -C1 for more details on baselines.
IV-A3 Hyperparameters
As in [40, 26], we implement the unconditional score model architecture with NCSN++M[59, 25], which is a convolution-based neural network operating in the complex STFT domain. NCSN++M is also used as the base architecture for PSE, SGMSE+ and StoRM. Details on the architecture, training configuration, reverberation operator and diffusion hyperparameters can be found in appendices -C1., -C1 and -C1, respectively.
IV-A4 Instrumental metrics
For instrumental evaluation of the speech dereverberation performance, we use the intrusive Perceptual Evaluation of Speech Quality (PESQ) [75] and extended short-term objective intelligibility (ESTOI) [76] for assessment of speech quality and intelligibility respectively. We also use the non-intrusive DNS-MOS [77], a DNN-based mean opinion score (MOS) approximation following the ITU-T P.835 recommendation [78]. We also report SI-SDR [79], however it must be noted that such point-wise distance metrics do not represent well the performance of generative models, given the natural variability of the corresponding estimates.
IV-A5 Instrumental evaluation results
We display in Table I the dereverberation results for all blind methods, both supervised and unsupervised. Blind supervised approaches PSE, SGMSE+ and StoRM generally perform better than unsupervised methods as they benefit from supervision at training time. However, we can observe the limited generalization ability of supervised approaches on the VCTK-SimulatedReverb when reverberant conditions are not the same as those presented during training. Our method BUDDy, however, seamlessly adapts to changing acoustics since it was trained without supervision. This enables BUDDy to retain its performance from VCTK-RealReverb to VCTK-SimulatedReverb, where supervised methods like PSE lose up to 0.77 PESQ points in mismatched reverberant conditions.
When evaluating on the DDS-DAPS-RealRecorded dataset, as reported in Table II, the generalization gap between our method and the supervised baselines increases even further. On this benchmark, BUDDy outperforms the best supervised baselines SGMSE+ and StoRM on DNS-MOS and ESTOI and has similar PESQ scores. This experiment highlights several points. First, BUDDy is naturally robust to background noise although it is not present in its signal model, echoing the conclusions of our prior work [26]. Informal listening suggests that the background noise is left untouched by the algorithm, and therefore it can be easily dealt with after processing. Furthermore, BUDDy’s unconditional diffusion model trained on the anechoic VCTK dataset seems to generalize well to DAPS speech utterances. Finally, BUDDy also performs strong dereverberation when handling a realistic reverberant model, since the utterances in the DDS-DAPS-RealRecorded dataset are directly recorded in a reverberant room, and not produced by a convolution between an utterance and a measured RIR. On the contrary, the supervised baselines suffer from the multiple mismatches between the presented benchmark and the training conditions, be it with regard to background noise, different speech corpora or reverberant model. This strengthens the position of BUDDy as a robust unsupervised baseline versus top-performing supervised baselines.
Overall, BUDDy performs far better than all other blind unsupervised baselines. For instance, BUDDy outperforms RVAE-EM by as much as 0.47 PESQ and 0.12 ESTOI points. Indeed, traditional unsupervised methods [7, 8] only draw limited benefits from their uninformed Gaussian prior on anechoic speech, while diffusion-based Saito et al. [34] and GibbsDDRM [35] seem to only marginally deviate from their WPE initialization. RVAE-EM [38] also obtains low instrumental scores, but informal listening suggested that its dereverberation abilities were superior to those of WPE.
TABLE II: Speech dereverberation results on DDS-DAPS-RealRecorded reverberant benchmark. Boldface numbers indicate best performance for supervised and unsupervised methods separately. Underlined numbers indicate best performance across all methods.
Method Unsup. DNS-MOS PESQ ESTOI
Reverberant - 2.35 $\pm$ 0.59 1.30 $\pm$ 0.16 0.63 $\pm$ 0.10
PSE 2.91 $\pm$ 0.55 1.65 $\pm$ 0.40 0.74 $\pm$ 0.14
SGMSE+M [24] 3.21 $\pm$ 0.52 2.14 $\pm$ 0.42 0.83 $\pm$ 0.09
StoRM [25] 3.48 $\pm$ 0.45 2.12 $\pm$ 0.50 0.82 $\pm$ 0.12
WPE [60] 2.64 $\pm$ 0.59 1.48 $\pm$ 0.29 0.70 $\pm$ 0.11
RVAE-EM [38] 2.68 $\pm$ 0.55 1.59 $\pm$ 0.33 0.71 $\pm$ 0.11
BUDDy (ours) 3.55$\pm$ 0.49 2.11 $\pm$ 0.47 0.86 $\pm$ 0.10
IV-A6 Ablation study
We conduct an ablation study to evaluate the impact of the projection step (15) introduced in the operator optimization (see Section III-A). We present the results in Table III and observe that, although the minimum-phase consistency projection has a theoretical justification as a mean to enhance the stability of the inverse RIR during optimization, its practical effect appears negligible, which can be due to a mismatch with the fact that real RIRs are generally mixed-phase filters [12]. However, we observe that the other operations in the projection step, i.e. STFT consistency, enforcement of the direct path, and speech magnitude constraint, are all instrumental in guiding BUDDy toward a solution with higher fidelity to clean speech, as measured by PESQ. We show DNS-MOS figures out of completeness. However, DNS-MOS variations are small across ablations and not indicative of fidelity to reference speech as DNS-MOS is not intrusive.
Additionally, we examine the effect of parameterizing the likelihood model with a $L^{2}$-distance on compressed spectrograms rather than on waveforms as in previous work [26]. To do so, we replace the cost function $C(\cdot,\cdot)$ from (9), which is based on compressed STFT representations, with a simpler waveform-domain $L^{2}$-distance, and we empirically pick the optimal corresponding scaling factor $\tilde{\zeta}$. The results clearly show the superiority of the proposed cost function (9) using compressed STFT representations.
IV-A7 Listening experiment
Instrumental metrics offer only limited insights into the performance of dereverberation algorithms [80]. We therefore conduct a listening experiment based on the MUSHRA recommendation [81] to assess the performance of BUDDy as perceived by human listeners. The test comprised 12 pages, featuring 6 reverberant speech utterances from the VCTK-RealReverb (matched) and the VCTK-SimulatedReverb (mismatched) sets. Participants were asked to rate the different stimuli with a single number representing overall quality, taking into account factors such as voice distortion, residual reverberation, and potential artifacts [80]. The test stimuli include our proposed method BUDDy, the unsupervised WPE [7] and RVAE-EM [38], as well as the supervised baselines PSE and SGMSE [24], Further details on the organization of the listening experiment are reported in Appendix -C1.
TABLE III: Ablation study on VCTK-RealReverb.
Method PESQ DNS-MOS
Reverberant 1.61 $\pm$ 0.37 3.14 $\pm$ 0.52
BUDDy 2.30 $\pm$ 0.53 3.76 $\pm$ 0.41
- Minimum-phase Consistency 2.30 $\pm$ 0.57 3.81 $\pm$ 0.40
  - RMS Power Constraint 2.22 $\pm$ 0.50 3.64 $\pm$ 0.50
    - Fixed Direct Path 2.10 $\pm$ 0.46 3.78 $\pm$ 0.44
      - STFT Consistency 1.96 $\pm$ 0.41 3.84 $\pm$ 0.39
$L^{2}$-Distance for $\mathcal{C}(\cdot,\cdot)$ 1.86 $\pm$ 0.47 3.36 $\pm$ 0.56
The results of the experiment are presented in Fig. 2. It can be observed that the unsupervised baselines WPE and RVAE-EM received low scores. Yet, RVAE-EM performs consistently better than WPE in this listening experiment, as opposed to what is suggested by instrumental metrics in Table I. In the matched test set (Fig. 2a), BUDDy obtained significantly lower scores than PSE and SGMSE+ ($p<0.001$ in a paired Welch test). However, in the mismatched set, PSE and SGMSE+ suffered a decrease in performance, losing up to $20$ points (out of $100$), while BUDDy mantained similar scores. In that case, there is no significant difference in performance between the three approaches ($p>0.1$), which closes the gap between BUDDy and the top-performing supervised baselines in this mismatched setting, highlighting the advantage provided by unsupervised learning.
Figure 2: Listening test results on VCTK-RealReverb and VCTK-SimulatedReverb. The boxplot shows first quartile, median, and third quartile.
IV-A8 Computational complexity
A current limitation of the proposed method lies in its high computational budget. Compared to classical diffusion approaches, such as SGMSE+ [24], our method requires several optimization iterations for the reverberation operator parameters at each reverse diffusion step, which increases the per-step inference time from 0.03s to 0.10s between SGMSE+ and BUDDy for each second of speech signal processed, as measured on a NVIDIA RTX A6000 GPU. Together with the larger number of reverse diffusion steps, the overall inference time is significantly higher than e.g. SGMSE+ or predictive approaches like PSE in Table I. Thus, while BUDDy is able to produce unprecedented high quality speech dereverberation without training on paired data, it cannot currently do so without an increase in computational complexity.
IV-B Singing Voice Dereverberation
We extend our evaluation benchmark to include the related task of singing voice dereverberation.
TABLE IV: Singing voice dereverberation results on NHSS dataset. Boldface numbers indicate best performance for supervised and unsupervised methods separately
Matched Mismatched
Method Unsup. $\ell^{1}$ STFT FAD $\ell^{1}$ STFT FAD
Reverberant - 1.98 $\pm$ 0.66 6.41 1.86 $\pm$ 0.56 5.65
PSE 1.56 $\pm$ 0.70 1.15 2.24 $\pm$ 0.78 1.79
SGMSE+ [24] 1.32 $\pm$ 0.51 0.82 1.37 $\pm$ 0.39 0.65
WPE [7] 2.02 $\pm$ 0.65 4.66 2.29 $\pm$ 0.66 5.74
Saito et al. [34] 1.95 $\pm$ 0.65 5.46 1.77 $\pm$ 0.52 4.90
BUDDy 1.90 $\pm$ 0.59 0.88 1.91 $\pm$ 0.50 0.60
Figure 3: Listening test results on singing voice datasets NHSS-RealReverb and NHSS-SimulatedReverb. The boxplot shows first quartile, median, and third quartile.
IV-B1 Data
We collect several publicly available singing voice datasets [82, 83, 84, 85, 86, 87]. These datasets feature over 94 h of studio-quality solo singing from a diverse array of singers and singing styles, spanning various languages. The majority of the recordings are in Chinese, followed by English, Japanese, and Korean. All datasets are down-sampled to 44.1 kHz. For testing, similar to [35], we use the sung part of NHSS [88, 89]. The NHSS dataset contains 100 English-language pop songs, 10 for each of the five male and five female singers recruited. We select a subset (90%) of the RIRs curated for the VCTK-based experiments, such that we only retain the RIRs whose original sample rate is at least 44.1 kHz. The resulting reverberant test set is referred to as NHSS-RealReverb. As for the speech voice experiments, we also prepare mismatched test set denoted as NHSS-SimulatedReverb using simulated RIRs sampled at 44.1 kHz.
IV-B2 Baselines
We evaluate the performance of BUDDy against two unsupervised baselines: WPE [7] and the unsupervised method from Saito et al. [34] which was originally designed for singing voice dereverberation. Additionally, we train supervised baselines following the same approach as in the speech dereverberation experiments described in Section IV-A2. Specifically, we compare BUDDy to PSE and SGMSE+, which share the same architectural design as BUDDy’s diffusion model, as detailed in Section IV-B3.
IV-B3 Hyperparameters
Similar to the speech dereverberation experiments, we adopt NCSN++ as the score model architecture. However, we adjust its hyperparameters to accommodate the higher sampling rate of 44.1 kHz. More details concerning the architecture and specific training configuration and inference hyperparameters are reported in Appendix -C2
IV-B4 Evaluation metrics
Objective metrics for evaluating singing voice restoration tasks are limited compared to those available for speech processing. Following [34], we use the $\ell^{1}$-distance in the magnitude STFT domain and a Fréchet Audio Distance (FAD) using a VGGish embedding [90]. However, these are only limited in interpretability and hardly relate to listening impression [91, 92]. Therefore, we complete this evaluation benchmark with a listening test with 10 participants, using a similar setup as reported in Section IV-A7. The test included 12 reverberant singing voice examples from the NHSS dataset, containing 6 reverberant singing voice examples from each of the matched and mismatched datasets. The instructions were identical to those reported in Appendix -C1.
IV-B5 Results
The results from the instrumental evaluation are reported in Table IV and those from the listening test in Fig. 3. The results show that BUDDy largely outperforms the unsupervised baselines and the supervised baseline PSE on FAD and subjective listening. In particular, in the mismatched test set, PSE fails to produce high-quality speech whereas BUDDy remains consistent with its good performance on the matched test set. BUDDy outperforms SGMSE+ on FAD in the mismatched setting, However, although it loses a few points in the mismatched scenario, SGMSE+ surpasses BUDDy by a small yet statistically significant margin on both test benchmarks in the listening experiment.
IV-C Robustness of RIR-Informed Methods
Informed dereverberation algorithms such as [5, 26] assume complete knowledge of the room acoustics as provided by the RIR $\mathbf{h}$. However, as pointed out in Section I, even if the RIR is perfectly known, single channel dereverberation is not trivial as RIRs are mixed-phase systems, such that causal and stable inverse filters do not exist [12]. We examine here the sensitivity of informed dereverberation methods in partially blind scenarios, i.e. when RIRs are known up to estimation errors.
IV-C1 Baselines
We include several informed dereverberation approaches for comparison. The baseline InfDerevDPS is inspired from previous work [26] and fully described in Section II-B. We also include the regularized inverse filtering method RIF+Post [5]. The second method RIF+Post [5] performs regularized inverse filtering in the Fourier domain, followed by traditional speech enhancement [93]. More details on these baselines can be found in Appendix -C1.
IV-C2 Synthetic RIR estimation errors
We first study the case where the oracle RIR is corrupted by Gaussian noise. The results displayed in Fig. 4 indicate that the performance of both the diffusion-based and the traditional method dwindles as the noise power increases. This suggests that informed methods have very limited robustness to errors in the provided RIR. This is a crippling drawback since obtaining a perfect sample-wise estimation of RIRs is an arduous problem, given their stochastic nature [11].
IV-C3 DNN-based RIR estimation errors
We now shift to a more realistic scenario where the RIR is blindly estimated from the reverberant speech by a DNN, since in practice RIR estimation errors are unlikely to be perfectly Gaussian-distributed. In particular, we employ FiNS [41], a state-of-the-art supervised RIR estimator which obtains RIR estimates based on the reverberant utterance (see Section IV-D1 for details). We compare in Table V results where the RIR is perfectly known (i.e. informed scenario) versus when it is estimated by FiNS (i.e. partially blind). The acoustic conditions in the considered evaluation set match those of the training set. Therefore, since FiNS was trained in a supervised fashion using paired reverberant/RIR data, it is expected to perform well on such conditions. The dereverberation performance of both InfDerevDPS and RIF+Post is very poor when the RIR is estimated with FiNS [41], as opposed to when the RIR is perfectly known. Yet, through informal listening, we notice that FiNS produces perceptually reasonable RIR estimates, which highlights the very limited robustness of informed methods when estimation errors, even imperceptible, affect the RIR knowledge. This all suggests that in blind cases the RIR should be jointly estimated with the anechoic speech, which is the paradigm followed by our method BUDDy.
01020301001111222233334444RIR Error SNR [dB]PESQ0102030100222233334444RIR Error SNR [dB]DNSMOSRIF+Post [5]InfDerevDPS [26]
Figure 4: Robustness of informed dereverberation approachess with respect to normally distributed errors in the RIR.
TABLE V: Dereverberation results on matched reverberant VCTK dataset. We indicate for each method in the table if it operates in a blind scenario.
Method Blind DNS-MOS PESQ
Reverberant - 3.14 $\pm$ 0.52 1.61 $\pm$ 0.37
RIF+Post [5] 3.41 $\pm$ 0.47 2.66 $\pm$ 0.40
InfDerevDPS [26] 3.91 $\pm$ 0.33 3.95 $\pm$ 0.42
FiNS/RIF+Post [41, 5] 2.18 $\pm$ 0.38 1.33 $\pm$ 0.19
FiNS/InfDerevDPS[41, 26] 2.19 $\pm$ 0.43 1.32 $\pm$ 0.18
BUDDy (ours) [40] 3.76 $\pm$ 0.41 2.30 $\pm$ 0.53
IV-D Room Impulse Response Estimation
Figure 5: RIR estimation metrics for each octave and full-band on the reverberant VCTK dataset. The violin plots show the distribution and the median. Lower is better. FiNS [41] is trained in a supervision fashion whereas BUDDy is unsupervised.
BUDDy is not only designed as a dereverberation algorithm but also functions as a blind unsupervised RIR estimator. We evaluate its performance for RIR estimation using the same speech model and data we employed for speech dereverberation in Section IV-A.
IV-D1 Baseline
We benchmark BUDDy against FiNS [41], a DNN-based approach trained to estimate time-domain RIRs directly from reverberant speech. FiNS comprises a 1D-convolutional encoder and a two-component decoder. The first decoder component models the late tail of the RIR by passing noise signals through a trainable filterbank containing several FIR filters. The second decoder component directly estimates the direct path and early reflections in the time-domain. In contrast to BUDDy, FiNS relies on supervised learning, thus requiring a paired dataset of reverberant speech and RIRs. We use an unofficial re-implementation [^2] https://github.com/kyungyunlee/fins and train the model on our VCTK-based reverberant speech dataset.
IV-D2 Evaluation metrics
Due to the highly ill-posed nature of the blind RIR estimation problem and the statistical nature of late reflections [11], we refrain from using element-wise distances, such as error-to-signal ratios, to evaluate the performance of RIR estimators. Instead, it is arguably more important to preserve the acoustic and perceptual properties of the reference RIR [94]. On the other hand, single metrics such as the full-band T60 reverberation time or clarity index C50 do not account for frequency-specific estimation errors. We therefore incorporate both full-band and subband reverberation time T60 and clarity index C50, with subbands spanning octaves. This enables to keep a high-level representation of the acoustical properties while allowing enough granularity on the spectral attributes of the RIR.
The reverberation time T60 is defined for a diffuse sound field as the time it takes for its energy decay curve (EDC) to decay by $60$ dB [1]. In order to avoid the effects of the noise floor, we calculate T60 as twice the time required for the EDC to decrease from $-5$ dB to $-35$ dB relative to the initial level, thereby eliminating the influence of the direct path. This measure is computed in each octave band separately. The octave clarity index C50 is the ratio (in dB) between the energy in the first $50$ ms and the energy in the remaining of the RIR, calculated in the corresponding octave band [1]. Consequently, we compute the absolute error between the T60 and C50 values calculated for each octave from the estimated RIR and those from the ground truth RIR.
IV-D3 Results
The results for both matched and mismatched test sets are plotted in Fig. 5. In the matched condition, FiNS and BUDDy achieve similar T60 error rates at low- and mid-range frequency bands, while BUDDy’s performance decreases at high frequencies (Fig. 5a). Our intuition is that the lower RIR estimation abilities of BUDDy at high frequencies can be related to the tendency of diffusion models to generate high-frequency components in the later stages of the reverse diffusion process [95]. Consequently, there is less information available for optimizing the RIR parameters in this range when beginning sampling, negatively affecting parameter convergence. A similar trend is observed for the C50 error in Fig. 5c. Furthermore, BUDDy generally achieves lower C50 error than FiNS in the mid-frequency range, where most of the speech content lies.
In the mismatched setting, FiNS struggles to generalize because of its supervised training setup. As a result, BUDDy outperforms FiNS in both T60 and C50 error at low and mid-frequency bands (Figs. 5b and 5d). At higher frequencies, BUDDy’s T60 estimation performance still remains slightly inferior to FiNS, though the gap is noticeably smaller than in the matched setting. Regarding C50, BUDDy outperforms FiNS in all frequency bands except for the highest 4-kHz band. This increased relative performance of BUDDy compared to FiNS highlights the benefits of leveraging unsupervised training for RIR estimation in variable acoustic conditions.
V Conclusion
In this paper, we presented an unsupervised method that simultaneously performs blind dereverberation and RIR estimation using diffusion models. Our results highlight the importance of joint speech and RIR estimation in contrast to plugging estimated RIRs into informed dereverberation methods. The proposed method, BUDDy, yields state-of-the-art performance among unsupervised approaches for blind speech and singing voice dereverberation, outperforming both traditional and DNN-based methods. Unlike blind supervised methods, which often struggle with generalization to unseen acoustic conditions, our unsupervised approach naturally overcomes this limitation due to its ability to adapt the reverberation operator to a broad range of RIRs. This holds as well for RIR estimation, as we show that the RIR estimation performance of BUDDy surpasses that of a state-of-the-art supervised DNN-based technique in mismatched acoustic conditions while being on par in a matched setting.
Acknowledgments
We would like to thank Koichi Saito, Fumikuri Yohena, and Kohei Yatabe for providing us with code and guidance through their methods. Many thanks to Julius Richter and Till Svajda for their advice on retraining SGMSE+ for singing voice dereverberation.